# The steep redshift evolution of the hierarchical binary black hole merger rate may cause the $z$-$\chi_{\rm eff}$ correlation

### v2

@@ Abstract
There is growing evidence from gravitational-wave observations that some merging black holes are created from previous mergers. Using the prediction that these hierarchically-merged black holes have dimensionless spin magnitudes of $\chi\approx 0.69$, we identify a subpopulation in the gravitational-wave data consistent with a hierarchical-merger origin in dense star clusters. This subpopulation’s primary mass distribution peaks at $16.7^{+3.1}_{-4.6}\,\mathrm{M}_{\odot}$, which is approximately twice as large as its secondary mass distribution’s mode ($8.0^{+29.7}_{-2.3}\,\mathrm{M}_{\odot}$), and its spin tilt distribution is consistent with isotropy. Our inferred secondary mass distributions imply that isolated binary evolution may still be needed to explain the entirety of the $9\,\mathrm{M}_{\odot}$ peak. Surprisingly, we find that the rate of hierarchical mergers may evolve more steeply with redshift than the rest of the population ($98.5\%$ credibility): the fraction of all binary black holes that are hierarchically formed at $z=0.1$ is $0.05^{+0.05}_{-0.04}$, compared to $0.17^{+0.13}_{-0.12}$ at $z=1$. This provides an explanation for the previously-discovered broadening of the effective spin distribution with redshift. Our results have implications for star cluster formation histories, as they suggest the potential existence of a high-redshift population of massive, compact clusters.
@@ II Phenomenological population model
Using a hierarchical Bayesian analysis (Loredo, 2004; Thrane and Talbot, 2019; Vitale et al., 2022b), we fit this model to all BBHs in GWTC-4, which includes BBHs in all previous catalogs. We additionally include GW241110 and GW241011 (Abac et al., 2025a), two events that have been recently published by the LVK because their spins and mass ratios are consistent with a hierarchical merger origin. However, excluding these events does not noticeably change our findings. We omit GW231123  (Abac et al., 2025b) as its large secondary mass (above $100\,M_{\odot}$) appears to be an outlier with respect to our inferred secondary mass distributions. Including this event does not affect our redshift or spin inferences. Appendix B further discusses the effects of GW231123 on our inferred mass distributions.
@@
We find strong evidence for a subpopulation of $\chi_{1}=0.69$ systems. The fraction of systems in this subpopulation, $\xi$, is larger than $0.02$ at the $99.7\%$ level, representing a Bayes factor of $20.3$. This subpopulation has spin orientation, mass ratio, and component mass distributions characteristic of hierarchical mergers in star clusters (see Appendices C.1, D, and Section V, respectively).
Marginalized over the range $z=[0,2]$, our inferred value of $\xi=0.23^{+0.18}_{-0.17}$ is roughly consistent with previous studies that also identify a 2G+1G-like subpopulation that makes up roughly $12-20\%$ of the population [^1] Antonini et al. (2025) estimate that $20\%$ of all mergers are 2G+1G, and Tong et al. (2025a) estimate the rate of 2G+1G mergers to be $2.2^{+1.9}_{-1.2}\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, which corresponds to $\approx 12\%$ of the total BBH merger rate inferred by The LIGO Scientific Collaboration et al. (2025), albeit with a different population model. Pierra et al. (2024) find that $\approx 2$$5\%$ of the overall population belong to a $\chi\approx 0.7$ subpopulation, which is lower than that inferred in this work, potentially caused by differing inferred redshift distributions. Additionally, Kimball et al. (2021) find $\mathcal{R}_{\rm 2G+1G}/\mathcal{R}_{\rm 1G+1G}<0.12$ when analyzing GWTC-2, and our posterior on this value peaks at $0.23$, putting these two studies performed on different datasets in mild tension. (e.g. Kimball et al., 2021; Wang et al., 2022; Mould et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tong et al., 2025b, a; Berti et al., 2025; Vijaykumar et al., 2026). The fact that similar fractions of 2G+1G systems are inferred – regardless of what compact binary parameter (i.e. effective spin, mass ratio, or primary spin magnitudes) is used to identify them – strengthens the evidence that these independently-discovered subpopulations are indeed the same as one another and represent 2G+1G mergers.
@@ IV The hierarchical merger rate evolves faster with redshift than the rest of the population
The merger rate of the $\chi_{1}\approx 0.7$ subpopulation may evolve more steeply than that of the main population. Figure 1 shows the merger rate evolution of both subpopulations as a function of redshift, as well as the posteriors on the power law index of each population’s redshift distributions, $\kappa_{\rm main}$ and $\kappa_{\rm sub}$. The power law index of the total population, as inferred by The LIGO Scientific Collaboration et al. (2025), is $3.2^{+0.94}_{-1.00}$, which lies between the values we infer for our main and subpopulations. The posterior on $\kappa_{\rm main}$ is less than that on $\kappa_{\rm sub}$ at $98.5\%$ credibility. There are therefore hints, but not yet definitive evidence, that hierarchical mergers represent a higher fraction of the binary black hole population at $z=1$ compared to $z=0$. If it is a real feature of the population, this result can explain the previously-observed broadening of the effective spin distribution with redshift (Biscoveanu et al., 2022), as we discuss in Section VI.
Hierarchical mergers necessarily proceed the first 1G mergers in a cluster, and therefore occur at similar or slightly later times in a clusters’ history than 1G mergers do. This expectation appears to be in contradiction with the results shown in Figure 1, implying either that our subpopulation is not a set of 2G+1G mergers (in contrast to our spin and mass ratio results), or that a more complex picture of star cluster contributions to the BBH merger rate is necessary to explain the data. Assuming a cluster mass function that scales inversely with the square of the cluster mass, the metallicity-specific star formation history from Madau and Fragos (2017), and the catalog of Cluster Monte Carlo simulations from Kremer et al. (2020), we infer the fraction of all BBHs that originated in star clusters to be $0.32^{+0.35}_{-0.23}$ at $z=0.1$ and $1.23^{+0.93}_{-0.84}$ at $z=1$. This directly follows from our measured fraction of 2G+1G mergers in the total population at different redshifts, which is $0.05^{+0.05}_{-0.04}$ at $z=0.1$ and $0.17^{+0.13}_{-0.12}$ at $z=1$. Thus, if our assumptions about the redshift-dependent mass function of clusters are accurate, contributions from other formation scenarios may be necessary to explain the merger rate at low redshift, but clusters can explain the full merger rate at high redshift. Additionally, the fact that the fraction of cluster-origin BBHs may be larger than unity at $z=1$ means that denser clusters at high redshift may be required to explain our inferred fraction of 2G+1G mergers. Theoretical uncertainties on cluster properties as a function of redshift remain uncertain, and future work will aim to constrain these with the subpopulation’s inferred redshift distribution (e.g. Fishbach and Fragione, 2023). We further discuss the potential astrophysical implications of our inferred redshift distributions in Section VII.
@@ V Mass distributions
Our inferred component mass distributions for each subpopulation are shown in Fig. 2. As in Farah et al. (2024), we display the underlying component mass distributions (i.e. after the pairing function has been factored out), but we provide marginalized component mass distributions in Appendix F. The hierarchical subpopulation’s primary mass distribution (representing 2G black holes) has a mode at $16.7^{+3.1}_{-4.6}\,\mathrm{M}_{\odot}$, which may be approximately twice as large as the mode of its secondary mass distribution ($8.0^{+29.7}_{-2.3}\,\mathrm{M}_{\odot}$). This is consistent with a 2G+1G interpretation of the subpopulation.
Additionally, the subpopulation contributes the majority of the merger rate at high primary masses ($\gtrsim 60\,\mathrm{M}_{\odot}$), as shown by the agreement between the orange (total) and green (subpopulation) bands in that region. The subpopulation’s secondary mass distribution does not extend to high masses, whereas the main population’s primary and secondary mass distributions appear to be similar. Even though their maximum masses are allowed to differ, the primary and secondary mass distributions of the main (1G+1G) population both exhibit negligible merger rates above $\approx 46\,\mathrm{M}_{\odot}$ (see Fig. 12 in Appendix F). Taken together, these results further support the hypothesis proposed in Tong et al. (2025b) that the pair-instability mass gap is polluted primarily by 2G+1G mergers, and not by 2G+2G mergers. However, it is possible that “second-generation”+“second-generation” mergers are present but not yet distinguishable from the rest of the population, and we cannot rule out their presence.
@@ VII Discussion
We find that the rate of hierarchical mergers likely evolves more steeply with redshift than the rest of the population. This would imply that 2G+1G mergers occur at earlier times than many of the the 1G+1G mergers observable in current gravitational wave data. This is potentially surprising, as hierarchical mergers cannot occur before “second-generation” black holes are made through “first-generation”+“first-generation” mergers. Ye and Fishbach (2024) shows that if the cluster mass function does not evolve with redshift, no significant difference is expected between the redshift evolution of 1G+1G versus hierarchical mergers (except at very early times before any “second-generation” black holes have formed). Our results that hierarchical mergers tend to occur at higher redshifts than “first-generation”+“first-generation” mergers, and that the fraction of hierarchical mergers may be as high as $0.17^{+0.13}_{-0.12}$ at $z=1$, may therefore imply that the cluster mass function does in fact evolve with redshift, and that hierarchical mergers originate in star clusters that are denser, more massive, and at higher redshift than the lower-redshift population that hosts the rest of the observed mergers (e.g. Mai et al., 2025). This interpretation is consistent with recent near-infrared observations, which suggest that high-redshift clusters are formed with higher densities, masses, and occurence rates than the low-redshift clusters against which many cluster simulation codes are calibrated (Vanzella et al., 2022; Mowla et al., 2022; Adamo et al., 2024; Mowla et al., 2024; Mayer et al., 2025).
@@
The authors gratefully acknowledge helpful conversations with Hui Tong, Daniel Holz, Utkarsh Mali, Mike Zevin, Sharan Banajiri, and Thomas Dent. We extend special thanks to Bart Ripperda for granting us access to the Bee workstation at CITA, where the computations for this work were performed. AV acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) (funding reference number 568580). MF acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) under grant RGPIN-2023-05511, the Alfred P. Sloan Foundation, and the Ontario Early Researcher Award. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). This research has made use of the Astrophysics Data System, funded by NASA under Cooperative Agreement 80NSSC21M00561.
@@ Appendix B GW231123
Including GW231123 does not impact our redshift or spin results. However, it does impact our inference of the secondary mass distribution. When included, the event appears to be fully absorbed in the subpopulation, and causes the maximum secondary mass of the subpopulation to shift to large values that are inconsistent with those inferred when GW231123 is excluded. We therefore deem GW231123 an outlier in $m_{2}$ and exclude it from the analysis in the main text of this paper. One possible interpretation of GW231123’s inconsistency with the inferred secondary mass distribution of the subpopulation is that it is not a 2G+1G merger. However, we refrain from making definitive or quantitative statements on this event.
@@ C.1 Spin Orientations
The main population, however, has support for both isotropic and aligned-spin contributions, consistent with findings in The LIGO Scientific Collaboration et al. (2025). Aligned-spin systems nominally constitute $52^{+27}_{-25}\%$ of the main population, as indicated by the hyperposterior on the fraction of main-population systems contained within a Gaussian distribution centered at unity. While it is possible to interpret this as the percentage of isolated binary black hole mergers, the hyperparameter governing the fraction of systems in the Gaussian component is an over-estimate of the the fraction of aligned-spin systems, given that broad Gaussian distributions emulate uniform distributions.
@@ C.2 Spin Magnitudes
The inferred spin magnitude distributions for our main and subpopulation are shown in Fig. 6. Remember that p($\chi_{1}$) for the subpopulation is a truncated Gaussian distribution with fixed mean and standard deviation, $\mu^{\chi_{1}}_{S}=0.69$, $\sigma_{S}^{\chi_{1}}=0.1$. All uncertainty in that distribution is therefore due to the hyperposterior on $\xi$, the mixture fraction between the main and subpopulation. The secondary spin magnitude distribution is fit to the data, and we find that it may be peaked between $0.08$ and $0.21$, indicating a preference for small but nonzero spin in the 1G components. However, the hyperposterior is broad and we are unable to draw definitive conclusions from current data. Furthermore, the spin magnitude hyperprior – shown in dashed lines in Fig. 6 – resembles the hyperposterior, so we caution against drawing conclusions from the secondary spin magnitude distribution of the subpopulation.
We assume that the two components of the main population’s systems are independent and identically distributed, and find them to be peaked strongly at $\mu_{M}^{\chi}=0.12^{+.04}_{-.05}$, again indicating a preference for small but nonzero spins, and consistent with The LIGO Scientific Collaboration et al. (2025).

### v3

@@ Abstract
There is growing evidence from gravitational-wave observations that some merging black holes are created from previous mergers. Using the prediction that these hierarchically-merged black holes have dimensionless spin magnitudes of $\chi\approx 0.69$, we identify a subpopulation in the gravitational-wave data consistent with a hierarchical-merger origin in dense star clusters. This subpopulation’s primary mass distribution peaks at $17.0^{+18.3}_{-4.4}\,\mathrm{M}_{\odot}$, which is approximately twice as large as its secondary mass distribution’s mode ($10.5^{+29.7}_{-4.7}\,\mathrm{M}_{\odot}$), and its spin tilt distribution is consistent with isotropy. Our inferred secondary mass distributions imply that isolated binary evolution may still be needed to explain the entirety of the $9\,\mathrm{M}_{\odot}$ peak. Surprisingly, we find that the rate of hierarchical mergers may evolve more steeply with redshift than the rest of the population ($98.0\%$ credibility): the fraction of all binary black holes that are hierarchically formed at $z=0.1$ is $0.03^{+0.05}_{-0.02}$, compared to $0.09^{+0.11}_{-0.07}$ at $z=1$. This provides an explanation for the previously-discovered broadening of the effective spin distribution with redshift. Our results have implications for star cluster formation histories, as they suggest the potential existence of a high-redshift population of massive, compact clusters.
@@ II Phenomenological population model
Using a hierarchical Bayesian analysis (Loredo, 2004; Thrane and Talbot, 2019; Vitale et al., 2022b), we fit this model to all BBHs in GWTC-4, which includes BBHs in all previous catalogs. We additionally include GW241110 and GW241011 (Abac et al., 2025a), two events that have been recently published by the LVK because their spins and mass ratios are consistent with a hierarchical merger origin. However, excluding these events does not noticeably change our findings. We omit GW231123  (Abac et al., 2025b) as its large secondary mass (above $100\,M_{\odot}$) appears to be an outlier with respect to our inferred secondary mass distributions. Including this event does not qualitatively affect our redshift or spin inferences, though it slightly reduces the statistical significance of the result shown in Fig. 1, and it visibly changes the subpopulation’s secondary mass distribution. Appendix B further discusses the effects of GW231123 on our inferred mass distributions.
@@
We find strong evidence for a subpopulation of $\chi_{1}=0.69$ systems. The fraction of systems in this subpopulation, $\xi$, is larger than $0.01$ at the $99.9\%$ level, representing a Bayes factor of $49.3$. This subpopulation has spin orientation, mass ratio, and component mass distributions characteristic of hierarchical mergers in star clusters (see Appendices C.1, D, and Section V, respectively).
Marginalized over the range $z=[0,2]$, our inferred value of $\xi=0.13^{+0.16}_{-0.10}$ is roughly consistent with previous studies that also identify a 2G+1G-like subpopulation that makes up roughly $12-20\%$ of the population [^1] Antonini et al. (2025) estimate that $20\%$ of all mergers are 2G+1G, and Tong et al. (2025a) estimate the rate of 2G+1G mergers to be $2.2^{+1.9}_{-1.2}\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, which corresponds to $\approx 12\%$ of the total BBH merger rate inferred by The LIGO Scientific Collaboration et al. (2025), albeit with a different population model. Pierra et al. (2024) find that $\approx 2$$5\%$ of the overall population belong to a $\chi\approx 0.7$ subpopulation, which is lower than that inferred in this work, potentially caused by differing inferred redshift distributions. Additionally, Kimball et al. (2021) find $\mathcal{R}_{\rm 2G+1G}/\mathcal{R}_{\rm 1G+1G}<0.12$ when analyzing GWTC-2, which is in agreement with our posterior on this value, which peaks at $0.07$. (e.g. Kimball et al., 2021; Wang et al., 2022; Mould et al., 2022; Pierra et al., 2024; Antonini et al., 2025; Tong et al., 2025b, a; Berti et al., 2025; Vijaykumar et al., 2026). The fact that similar fractions of 2G+1G systems are inferred – regardless of what compact binary parameter (i.e. effective spin, mass ratio, or primary spin magnitudes) is used to identify them – strengthens the evidence that these independently-discovered subpopulations are indeed the same as one another and represent 2G+1G mergers.
@@ IV The hierarchical merger rate evolves faster with redshift than the rest of the population
The merger rate of the $\chi_{1}\approx 0.7$ subpopulation may evolve more steeply than that of the main population. Figure 1 shows the merger rate evolution of both subpopulations as a function of redshift, as well as the posteriors on the power law index of each population’s redshift distributions, $\kappa_{\rm main}$ and $\kappa_{\rm sub}$. The power law index of the total population, as inferred by The LIGO Scientific Collaboration et al. (2025), is $3.2^{+0.94}_{-1.00}$, which lies between the values we infer for our main and subpopulations. The posterior on $\kappa_{\rm main}$ is less than that on $\kappa_{\rm sub}$ at $98.0\%$ credibility. There are therefore hints, but not yet definitive evidence, that hierarchical mergers represent a higher fraction of the binary black hole population at $z=1$ compared to $z=0$. If it is a real feature of the population, this result can explain the previously-observed broadening of the effective spin distribution with redshift (Biscoveanu et al., 2022), as we discuss in Section VI.
Hierarchical mergers necessarily proceed the first 1G mergers in a cluster, and therefore occur at similar or slightly later times in a clusters’ history than 1G mergers do. This expectation appears to be in contradiction with the results shown in Figure 1, implying either that our subpopulation is not a set of 2G+1G mergers (in contrast to our spin and mass ratio results), or that a more complex picture of star cluster contributions to the BBH merger rate is necessary to explain the data. Assuming a cluster mass function that scales inversely with the square of the cluster mass, the metallicity-specific star formation history from Madau and Fragos (2017), and the catalog of Cluster Monte Carlo simulations from Kremer et al. (2020), we infer the fraction of all BBHs that originated in star clusters to be $0.18^{+0.30}_{-0.14}$ at $z=0.1$ and $0.74^{+0.89}_{-0.55}$ at $z=1$. This directly follows from our measured fraction of 2G+1G mergers in the total population at different redshifts, which is $0.03^{+0.05}_{-0.02}$ at $z=0.1$ and $0.09^{+0.11}_{-0.07}$ at $z=1$.
Thus, if our assumptions about the redshift-dependent mass function of clusters are accurate, contributions from other formation scenarios are necessary to explain the merger rate at low redshift, but clusters can explain the full merger rate at high redshift. Additionally, the fact that the fraction of cluster-origin BBHs at $z=1$ has support above unity means that denser clusters at high redshift may be required to explain our inferred fraction of 2G+1G mergers. Theoretical uncertainties on cluster properties as a function of redshift remain uncertain, and future work will aim to constrain these with the subpopulation’s inferred redshift distribution (e.g. Fishbach and Fragione, 2023). We further discuss the potential astrophysical implications of our inferred redshift distributions in Section VII.
@@ V Mass distributions
Our inferred component mass distributions for each subpopulation are shown in Fig. 2. As in Farah et al. (2024), we display the underlying component mass distributions (i.e. after the pairing function has been factored out), but we provide marginalized component mass distributions in Appendix F. The hierarchical subpopulation’s primary mass distribution (representing 2G black holes) has a mode at $17.0^{+18.3}_{-4.4}\,\mathrm{M}_{\odot}$, which may be approximately twice as large as the mode of its secondary mass distribution ($10.5^{+29.7}_{-4.7}\,\mathrm{M}_{\odot}$). This is consistent with a 2G+1G interpretation of the subpopulation.
Additionally, the subpopulation contributes the majority of the merger rate at high primary masses ($\gtrsim 60\,\mathrm{M}_{\odot}$), as shown by the agreement between the orange (total) and green (subpopulation) bands in that region. The subpopulation’s secondary mass distribution does not extend to high masses, whereas the main population’s primary and secondary mass distributions appear to be similar. Even though their maximum masses are allowed to differ, the primary and secondary mass distributions of the main (1G+1G) population both exhibit negligible merger rates above $\approx 46\,\mathrm{M}_{\odot}$ (see Fig. 12 in Appendix F). Taken together, these results further support the hypothesis proposed in Tong et al. (2025b) that the pair-instability mass gap is polluted primarily by 2G+1G mergers rather than by 2G+2G mergers. However, it is possible that “second-generation”+“second-generation” mergers are present but not yet distinguishable from the rest of the population, and we cannot rule out their presence.
@@ VII Discussion
We find that the rate of hierarchical mergers likely evolves more steeply with redshift than the rest of the population. This would imply that 2G+1G mergers occur at earlier times than many of the the 1G+1G mergers observable in current gravitational wave data. This is potentially surprising, as hierarchical mergers cannot occur before “second-generation” black holes are made through “first-generation”+“first-generation” mergers. Ye and Fishbach (2024) shows that if the cluster mass function does not evolve with redshift, no significant difference is expected between the redshift evolution of 1G+1G versus hierarchical mergers (except at very early times before any “second-generation” black holes have formed). Our results that hierarchical mergers tend to occur at higher redshifts than “first-generation”+“first-generation” mergers, and that the fraction of hierarchical mergers may be as high as $0.09^{+0.11}_{-0.07}$ at $z=1$, may therefore imply that the cluster mass function does in fact evolve with redshift, and that hierarchical mergers originate in star clusters that are denser, more massive, and at higher redshift than the lower-redshift population that hosts the rest of the observed mergers (e.g. Mai et al., 2025). This interpretation is consistent with recent near-infrared observations, which suggest that high-redshift clusters are formed with higher densities, masses, and occurence rates than the low-redshift clusters against which many cluster simulation codes are calibrated (Vanzella et al., 2022; Mowla et al., 2022; Adamo et al., 2024; Mowla et al., 2024; Mayer et al., 2025).
@@
The authors gratefully acknowledge helpful conversations with Hui Tong, Daniel Holz, Utkarsh Mali, Mike Zevin, Sharan Banagiri, and Thomas Dent. We extend special thanks to Bart Ripperda for granting us access to the Bee workstation at CITA, where the computations for this work were performed. AV acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) (funding reference number 568580). MF acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) under grant RGPIN-2023-05511, the Alfred P. Sloan Foundation, and the Ontario Early Researcher Award. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). This research has made use of the Astrophysics Data System, funded by NASA under Cooperative Agreement 80NSSC21M00561.
@@ Appendix B GW231123
Including GW231123 does not qualitatively impact our redshift or spin results, but it does lessen the statistical significance of the finding that the subpopulation has a steeper redshift evolution than the main population. Because GW231123 has high primary spins, it appears to be fully absorbed in the subpopulation. It’s redshift is also relatively low ($z\approx 0.4$), which causes the subpopulation’s inferred redshift distribution to flatten. Thus, we find $\kappa_{M}<\kappa_{S}$ to $88\%$ when GW231123 is included in the analysis, compared to 98% when it is excluded.
Furthermore, including GW231123 in the analysis impacts our inferred secondary mass distribution. When included, the event causes the maximum secondary mass of the subpopulation to shift to large values that are inconsistent with those inferred when GW231123 is excluded. We therefore deem GW231123 an outlier in $m_{2}$ and exclude it from the analysis in the main text of this paper. One possible interpretation of GW231123’s inconsistency with the inferred secondary mass distribution of the subpopulation is that it is not a 2G+1G merger. However, we refrain from making definitive or quantitative statements on this event.
@@ C.1 Spin Orientations
The main population, however, has support for both isotropic and aligned-spin contributions, consistent with findings in The LIGO Scientific Collaboration et al. (2025). Aligned-spin systems nominally constitute $62^{+24}_{-25}\%$ of the main population, as indicated by the hyperposterior on the fraction of main-population systems contained within a Gaussian distribution centered at unity. While it is possible to interpret this as the percentage of isolated binary black hole mergers in the 1G+1G population, the hyperparameter governing the fraction of systems in the Gaussian component is typically an over-estimate of the fraction of aligned-spin systems, given that broad Gaussian distributions emulate uniform distributions.
@@ C.2 Spin Magnitudes
The inferred spin magnitude distributions for our main and subpopulation are shown in Fig. 6. Remember that p($\chi_{1}$) for the subpopulation is a truncated Gaussian distribution with fixed mean and standard deviation, $\mu^{\chi_{1}}_{S}=0.69$, $\sigma_{S}^{\chi_{1}}=0.1$. All uncertainty in that distribution is therefore due to the hyperposterior on $\xi$, the mixture fraction between the main and subpopulation. The secondary spin magnitude distribution is fit to the data, and we find that it may be peaked between $0.02$ and $0.23$, indicating a preference for small but not necessarily zero spin in the 1G components. However, the hyperposterior is broad and the spin magnitude hyperprior – shown in dashed lines in Fig. 6 – resembles the hyperposterior for the subpopulation, so we further caution against drawing conclusions from the secondary spin magnitude distribution of the subpopulation.
We assume that the two components of the main population’s systems are independent and identically distributed, and find them to be peaked strongly at $\mu_{M}^{\chi}=0.12^{+.06}_{-.06}$, again indicating a preference for small but nonzero spins, and consistent with The LIGO Scientific Collaboration et al. (2025).